OpenStudy (anonymous):

Differentiate: x^3(sin(5x)^2 Help please

OpenStudy (anonymous):

you have to use product rule and chain rule

OpenStudy (anonymous):

\[x ^{3} \times \sin(5x)^{2}\]

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

yes got that bit :)

OpenStudy (anonymous):

please stay with me here

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

lets first find dy/dx of sin(5x)^2

OpenStudy (anonymous):

can you work on that?

OpenStudy (anonymous):

2sin(5x) * 5?

OpenStudy (anonymous):

sorry cos

OpenStudy (anonymous):

nooooo wrong

OpenStudy (anonymous):

\[dy/dx =\sin(5x)^2= 2\sin(5x)\cos(5x)\]

OpenStudy (anonymous):

now find dy/dx of x^3

OpenStudy (anonymous):

3x^2

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

now apply product rule

OpenStudy (anonymous):

\[U.dv/dx + V.du/dx\]

OpenStudy (anonymous):

let U=x^3 and V=sin(5x)^2

OpenStudy (anonymous):

now apply the formular

OpenStudy (anonymous):

dv/dx=\[2\sin(5x)\cos(5x)\]

OpenStudy (anonymous):

and du/dx=\[3x^2\]

OpenStudy (anonymous):

now our result will be

OpenStudy (anonymous):

\[2x^3\cos(5x)\sin(5x) + 3x^2\sin(5x)^{2}\]

OpenStudy (anonymous):

whic is same as \[x^2(2xcos(5x)\sin(5x)+3\sin(5x)^2)\]

OpenStudy (anonymous):

and that is the result

OpenStudy (anonymous):

excellent thank you

OpenStudy (anonymous):

thanks to openstudy

OpenStudy (anonymous):

you didn't differentiate the 5x?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ok. did a mistake there

OpenStudy (anonymous):

\[dy/dx = [x^3 *(d/dx) (\sin^2(5x))] + [(\sin^2(5x)) * (d/dx) x^3] \]

OpenStudy (anonymous):

\[ [x^3 * 2\sin(5x) * (d/dx) \sin(5x)] + [(\sin^2(5x)) * 3x^2] \]

OpenStudy (anonymous):

look at this step \[ [x^3 * 2\sin(5x) * \cos(5x) * (d/dx) 5x] + [(\sin^2(5x)) * 3x^2] \]

OpenStudy (anonymous):

\[ [x^3 * 2\sin(5x) * \cos(5x) * 5] + [(\sin^2(5x)) * 3x^2] \]

OpenStudy (anonymous):

\[ [10x^3 * \sin(5x) * \cos(5x)] + [3x2 * \sin^2(5x)] \]

OpenStudy (anonymous):

thats the result

OpenStudy (anonymous):

please close the question

OpenStudy (irishboy123):

is this \[\frac{d (x^3\sin(5x)^2)}{dx} = \frac{d (x^3\sin(25x^2))}{dx}\] or \[\frac{d (x^3\sin^2(5x))}{dx}\] looks like the former, has been done as the latter...